OpenStudy (anonymous):

Find the average value of the function on the given interval f(x)=cos x; interval \[[\frac{ -\pi }{ 2 },\frac{ \pi }{2 }]\]

OpenStudy (anonymous):

well the average value of f from -pi/2 to pi/2 is the integral \[\frac{ 1 }{ \Pi } \times \int\limits_{\frac{ -\Pi }{ 2 }}^{\frac{ \Pi }{ 2 }} \cos x dx = \frac{ 1 }{ \Pi } \times \sin x |_{\frac{ -\Pi }{ 2 }}^{\frac{ \Pi }{ 2 }}\]

OpenStudy (anonymous):

is it \[\frac{ 1 }{\pi}*(\sin \frac{ \pi }{ 2 }-\sin \frac{ -\pi }{ 2 })\] which would simplify down to \[\frac{ 2 }{ \pi } \]

OpenStudy (anonymous):

yes it's \[\frac{ 2 }{ π }\]

OpenStudy (anonymous):

nice and I found it properly correct? It wasn't just luck that it came out that way?

OpenStudy (anonymous):

yes, your result is accurate.

OpenStudy (anonymous):

thnx a bunch. well earned medal